OpenStudy (anonymous):

Perimeter is 55 cm, find length of each side....

OpenStudy (anonymous):

What shape?

OpenStudy (anonymous):



OpenStudy (anonymous):

rectange....trapezoid...?

OpenStudy (anonymous):

top: x+3 right: x^2-3x left: x+7 bottom: 3x-3

OpenStudy (anonymous):

quadralateral

OpenStudy (anonymous):

Okay! Perimeter is defined by all of the sides added together, right?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

So: \[55=(x+3)+(x^2-3)+(x+7)+(3x-3)\] 55, the perimeter is equal to all of the sides added together! We can combine like terms...

OpenStudy (anonymous):

So then you have \[55=x^2+5x+4\] I think.

OpenStudy (anonymous):

Can you solve for x?

OpenStudy (anonymous):

Oh no! I made a mistake. one moment.

OpenStudy (anonymous):

55=(x+3)+(x2−3x)+(x+7)+(3x−3)

OpenStudy (anonymous):

Which is then: x^2+2x+7

OpenStudy (anonymous):

Set it equal to 55 and solve.

OpenStudy (anonymous):

It says i need more than one answer, one for each side

OpenStudy (anonymous):

Each side of what?

OpenStudy (anonymous):

the quad

OpenStudy (anonymous):

Well if you solve for x, then you put x back into the equation for each side. Does that make sense?

OpenStudy (anonymous):

what is x? then i take that and put in each different equation?

OpenStudy (anonymous):

Yes. I can help you solve for x.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

So, you have 55=(x+3)+(x^2−3x)+(x+7)+(3x−3). This simplifies when you combine like terms to 55=x^2+2x+7. Can you solve it from that point by factoring or do you need more help?

OpenStudy (anonymous):

I dont know, what do i do with 55? i can only solve the other side

OpenStudy (anonymous):

Well, we can set the expression equal to zero: Move the 55 over to the right. So now, we have 0=x^2+2x-48. Do you see how I got that? Next, try to factor this; it works out nicely. Should give you two answers.

OpenStudy (anonymous):

Yes, and ill solve that, be right back :)

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

6 and -8?

OpenStudy (anonymous):

Yes! Okay, so we're talking about distance here. can we have a negative distance? Which solution do you think we should use?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yes. So, what are your dimensions if you plug in x=6?

OpenStudy (anonymous):

i got 9, 13, 15, and 18 total, are these correct?

OpenStudy (anonymous):

That's what I got! Good work!

OpenStudy (anonymous):

:D thanks

OpenStudy (anonymous):

:D You're very welcome. Have a good night.